OpenStudy (anonymous):

solve : 10x-4=5x+11+2x

OpenStudy (anonymous):

324

OpenStudy (anonymous):

10x - 4 = 7x +11 10x-7x = 11+4 3x = 15 x = 15/3 x = 5